# LargeAD: Large-Scale Cross-Sensor Data Pretraining for Autonomous Driving

### v1

@@ Abstract
Recent advancements in vision foundation models (VFMs) have revolutionized visual perception in 2D, yet their potential for 3D scene understanding, particularly in autonomous driving applications, remains underexplored. In this paper, we introduce LargeAD, a versatile and scalable framework designed for large-scale 3D pretraining across diverse real-world driving datasets. Our framework leverages VFMs to extract semantically rich superpixels from 2D images, which are aligned with LiDAR point clouds to generate high-quality contrastive samples. This alignment facilitates cross-modal representation learning, enhancing the semantic consistency between 2D and 3D data. We introduce several key innovations: i) VFM-driven superpixel generation for detailed semantic representation, ii) a VFM-assisted contrastive learning strategy to align multimodal features, iii) superpoint temporal consistency to maintain stable representations across time, and iv) multi-source data pretraining to generalize across various LiDAR configurations. Our approach delivers significant performance improvements over state-of-the-art methods in both linear probing and fine-tuning tasks for both LiDAR-based segmentation and object detection. Extensive experiments on eleven large-scale multi-modal datasets highlight our superior performance, demonstrating the adaptability, efficiency, and robustness in real-world autonomous driving scenarios. The project page can be accessible at: https://ldkong.com/LargeAD.
@@ 1 Introduction
The emergence of large language models (LLMs) [1, 2, 3, 4, 5] has transformed natural language processing, paving the way for similar breakthroughs in computer vision through vision foundation models (VFMs) such as SAM [6], X-Decoder [7], and SEEM [8]. These models have demonstrated remarkable capabilities in extracting rich pixel-level semantics from 2D images. However, extending these advancements to the 3D domain remains a largely unexplored frontier. As autonomous driving applications increasingly rely on 3D data from LiDAR sensors, there is a growing need to transfer the success of VFMs in 2D vision to 3D scene understanding [9, 10].
Accurate segmentation and detection of LiDAR point clouds are essential for safe autonomous driving and advanced driver-assistance systems [11, 12, 13, 14, 15]. Traditional LiDAR point cloud models often depend on large annotated datasets, which are costly and time-consuming to create [16, 17]. To alleviate this challenge, research has explored semi-supervised [18, 19] and weakly-supervised [17, 20] methods. However, these approaches suffer from limited generalizability, especially when faced with diverse sensor configurations, e.g., different LiDAR beam numbers, camera placements, sampling rates, and potential sensor corruptions [21, 22, 11, 23, 24, 25]. This limitation poses a significant challenge for real-world scalability.
@@
A core innovation in our framework is the use of VFMs to generate semantically enriched superpixels from camera images, which are then aligned with LiDAR data to construct high-quality contrastive samples (see Fig. 1). These semantic superpixels provide enhanced 2D-3D correspondences that capture object-level coherence, reducing the errors commonly associated with over-segmentation and “self-conflict” in contrastive learning [9]. This alignment significantly improves performance in downstream tasks, including 3D object detection and segmentation.
@@
Figure 1: Comparisons between i) the conventional image-to-LiDAR data pretraining frameworks [27, 28, 9] and ii) our proposed large-scale cross-sensor data pretraining (LargeAD). Our approach combines heterogeneous data sources for representation learning, which achieves superior robustness and scalability. Different from previous work, our framework encourages representation learning across different datasets, which largely enhances the generalizability.
@@
Figure 2: Illustration of image-to-LiDAR data pretraining using i) the heuristic SLIC algorithm [29] and ii) different vision foundation models (VFMs). Images in the first row are the superpixels generated by different methods, where each color represents one distinct segment. The LiDAR point clouds from the second row are the superpoints grouped by projecting superpixels to 3D using camera-LiDAR correspondence. The third row shows the linear probing results after data pretraining.
@@ 2.1 LiDAR-Based Scene Understanding
For autonomous vehicles, accurate and dense 3D perception is crucial for safe navigation [12, 30]. Researchers have developed various methods for point cloud segmentation, including those based on raw points [31, 32, 33, 34, 35], range views [36, 37, 38, 39, 40, 41, 42], bird’s eye views [43, 44], voxels [45, 46, 47, 48, 49], and multi-view fusion [50, 51, 52, 53, 54, 55]. Despite achieving significant advancements, these models typically rely on extensively annotated datasets, posing scalability issues [21]. To alleviate the annotation burden, recent studies have explored semi-supervised [18, 19, 56], weakly-supervised [17, 57, 58, 59, 20], and active learning [60, 61, 62] methods, as well as domain adaptation techniques [63, 64, 65, 66, 67, 68]. This work utilizes a self-supervised learning strategy, distilling knowledge from VFMs through camera-to-LiDAR associations, thereby eliminating the need for manual annotations during pretraining.
@@ 2.3 Representation Learning in 3D
3D self-supervised learning has its roots in image-based techniques and typically focuses on object-centric point clouds [73, 74, 75, 76, 77] or indoor scenes [78, 79, 80, 81, 82] using pretext tasks [83, 84, 85, 86, 87], contrastive learning [88, 89, 90, 91, 92, 93, 94], or mask modeling [95, 96, 97, 98]. These methods often lack the necessary scale and diversity for outdoor driving scenes [99, 100, 101]. Efforts such as PointContrast [102], DepthContrast [103], and SegContrast [104] have pioneered contrastive objectives for small-scale point clouds. Recently, Sautier et al. [27] introduced SLidR, the first method for image-to-LiDAR representation distillation in cross-modal self-supervised learning on large-scale point clouds. Mahmoud et al. [28] further refined this approach with semantically tolerant contrastive constraints and a class-balancing loss. SuperFlow [105] introduced a spatiotemporal consistency framework to efficiently capture dynamic cues across multiple timesteps. Our framework builds upon SLidR [27] by leveraging VFMs [6, 7, 8] to create a more effective cross-modal contrastive objective. We also introduce a superpoint temporal consistency regularization to enhance feature learning and robustness in diverse and dynamic real-world driving scenarios.
@@ 2.4 Multi-Dataset Utilization
Leveraging multiple datasets has emerged as a promising approach for improving the generalization of LiDAR-based models in autonomous driving [106]. Recent works like MDT3D [107] and MS3D++ [108] explored multi-source training for 3D object detection, while addressing challenges such as label space conflicts. Similarly, methods like COLA [109] and M3Net [110] utilized unified label spaces for semantic segmentation, demonstrating the advantages of multi-dataset learning. In this work, we extend these efforts by proposing a multi-source pretraining strategy that incorporates data from diverse LiDAR datasets, each with unique sensor setups and conditions [111, 112, 113]. By aligning 2D and 3D features across different sources, our approach improves cross-modal consistency and robustness. To the best of our knowledge, this is the first comprehensive study on multi-source pretraining for image-to-LiDAR representation learning, allowing our framework to generalize more effectively across varied driving environments and sensor configurations.
@@ 3 Image-to-LiDAR Data Pretraining
Figure 3: Overview of the VFM-driven image-to-LiDAR contrastive learning framework. Given a pair of LiDAR point cloud $\mathcal{P}^{t}$ and camera image $\mathcal{I}^{t}$ captured at timestamp $t$, along with another LiDAR point cloud $\mathcal{P}^{t+n}$ captured at timestamp $t+n$, we generate semantic superpixels using vision foundation models (VFMs). The corresponding superpoints are obtained by projecting image pixels onto the point cloud. Two key objectives are established: i) spatial contrastive learning between paired LiDAR and camera features, and ii) temporal consistency regularization between point segments from $\mathcal{P}^{t}$ and $\mathcal{P}^{t+n}$.
@@ 3.1 Problem Formulation
Let us define a point cloud $\mathcal{P}=\{\mathbf{p}_{i},\mathbf{e}_{i}\mid i=1,\ldots,N\}$ consisting of $N$ points collected by a LiDAR sensor. Each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ denotes the 3D coordinates, while $\mathbf{e}_{i}\in\mathbb{R}^{L}$ represents its feature embedding, such as intensity, elongation, etc. This work aims to transfer knowledge from a set of surround-view images =$\mathcal{I}=\{\mathbf{I}_{i}\mid i=1,\ldots,V\}$, captured by a total of $V$ synchronized RGB cameras, to the point cloud $\mathcal{P}$. Each image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ has a spatial resolution defined by height $H$ and width $W$.
Given that the LiDAR and camera sensors are assumed to be well-calibrated, each LiDAR point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ can be projected onto the image plane as a pixel $\hat{\mathbf{p}}_{i}=(u_{i},v_{i})$ using the following coordinate transformation:
@@ 3.3 Contrastive Learning Objective
Let $F_{\theta_{p}}:\mathbb{R}^{N\times(3+L)}\to\mathbb{R}^{N\times C}$ represent a LiDAR point cloud encoder with trainable parameters $\theta_{p}$, which processes a point cloud $\mathcal{P}$ and outputs a $C$–dimensional feature for each point. Additionally, let $G_{\theta_{i}}:\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times E}$ be an image encoder with parameters $\theta_{i}$, initialized from 2D self-supervised pretrained models. To compute the superpixel-driven contrastive loss, we construct trainable projection heads $H_{\omega_{p}}$ and $H_{\omega_{i}}$ which map the 3D point features and 2D image features into the same $D$-dimensional embedding space. The point projection head $H_{\omega_{p}}:\mathbb{R}^{N\times C}\to\mathbb{R}^{N\times D}$ is a linear layer followed by $\ell_{2}$-normalization. The image projection head $H_{\omega_{i}}:\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times E}\to\mathbb{R}^{H\times W\times D}$ consists of a convolution layer with a $1\times 1$ kernel, followed by a fixed bi-linear interpolation layer in the spatial dimension, and outputs with $\ell_{2}$-normalization.
The goal is to distill the 2D network’s knowledge into the 3D network, ensuring that each semantic superpoint feature closely correlates with its corresponding semantic superpixel feature. Specifically, the superpixels $\Phi_{\mathcal{S}}$ and superpoints $\Phi_{\mathcal{O}}$ are used to group pixel and point embedding features, respectively. An average pooling operation is applied to the grouped pixel and point embeddings to obtain superpixel embedding features $\mathbf{Q}\in\mathbb{R}^{M\times D}$ and superpoint embedding features $\mathbf{K}\in\mathbb{R}^{M\times D}$. The contrastive loss $\mathcal{L}^{\text{slic}}=\mathcal{L}\left(\mathbf{Q},\mathbf{K}\right)$ is then defined as follows:
[…]
where $\mathbf{q}_{i},\mathbf{k}_{j}\right>$ represents the scalar product between superpoint and superpixel embedding features, measuring their similarity. $\tau$ is a temperature parameter used to scale the similarity scores.
@@ 4.2 Semantic Spatial Consistency Learning
To implement this framework, we use the same trainable LiDAR point cloud encoder $F_{\theta_{p}}$ and frozen image encoder $G_{\theta_{i}}$ as described earlier, extracting features from the LiDAR point cloud and 2D images, respectively. For the contrastive loss, we employ the projection heads $H_{\omega_{p}}$ and $H_{\omega_{i}}$ from Sec. 3.2, projecting both point and image features into a shared $D$–dimensional embedding space. Unlike the low-level cues generated by SLIC, the superpixels produced by VFMs are enriched with semantic information, leading to more coherent and meaningful representations.
To compute the VFM-assisted contrastive loss, we apply average pooling to the pixel and point embeddings grouped by the superpixel set $\widehat{\Phi}_{\mathcal{S}}$ and the corresponding superpoint set $\widehat{\Phi}_{\mathcal{O}}$. This process yields superpixel embeddings $\widehat{\mathbf{Q}}\in\mathbb{R}^{M_{v}\times D}$ and superpoint embeddings $\widehat{\mathbf{K}}\in\mathbb{R}^{M_{v}\times D}$. The VFM-assisted contrastive loss $\mathcal{L}^{\text{vfm}}=\mathcal{L}\left(\widehat{\mathbf{Q}},\widehat{\mathbf{K}}\right)$ is formulated as follows:
@@ 4.3 Instance Superpoint Temporal Consistency
Superpoint Temporal Consistency. We leverage the clustered segments to compute the temporal consistency loss among related semantic superpoints [^1] Here, we assume $n=1$, that is, the current and the next LiDAR frames, without loss of generalizability.. Specifically, given a sampled temporal pair $\widetilde{\mathcal{P}}^{t}$ and $\widetilde{\mathcal{P}}^{t+1}$ and their corresponding segments $\widetilde{\mathcal{K}}^{t}$ and $\widetilde{\mathcal{K}}^{t+1}$, we compute the point-wise features ^$\widehat{\mathcal{F}}^{t}\in\mathbb{R}^{N\times D}$ and ^$\widehat{\mathcal{F}}^{t+1}\in\mathbb{R}^{N\times D}$ from the point projection head $H_{\omega_{p}}$. As for the target embedding, we split the point features ^$\widehat{\mathcal{F}}^{t}$ and ^$\widehat{\mathcal{F}}^{t+1}$ into $M_{k}$ groups by segments $\widetilde{\mathcal{K}}^{t}$ and $\widetilde{\mathcal{K}}^{t+1}$. Then, we apply an average pooling operation on ^$\widehat{\mathcal{F}}^{t+1}$ to get $M_{k}$ target mean feature vectors ^$\widehat{\mathcal{F}}^{t+1}=\{\widehat{\mathcal{F}}^{t+1}_{1},\widehat{\mathcal{F}}^{t+1}_{2},...,\widehat{\mathcal{F}}^{t+1}_{M_{k}}\}$, where ^$\widehat{\mathcal{F}}^{t+1}_{M_{k}}\in\mathbb{R}^{1\times D}$. Let the split point feature ^$\widehat{\mathcal{F}}^{t}$ be ^$\widehat{\mathcal{F}}^{t}=\{\widehat{\mathcal{F}}^{t}_{1},\widehat{\mathcal{F}}^{t}_{2},...,\widehat{\mathcal{F}}^{t}_{M_{k}}\}$, where ^$\widehat{\mathcal{F}}^{t}_{M_{k}}\in\mathbb{R}^{k\times D}$ and $k$ is the number of points in the corresponding segment. We compute the temporal consistency loss ${\mathcal{L}}^{t\to{t+1}}$ to minimize the differences between the point features in the current frame (timestamp $t$) and the corresponding segment mean features from the next frame (timestamp $t+1$) as ${\mathcal{L}}^{t\to{t+1}}$ as follows:
[…]
@@
Since the target embedding for all points within a segment in the current frame serves as the mean segment representation from the next frame, this loss will force points from a segment to converge to a mean representation while separating from other segments, implicitly clustering together points from the same instance. Fig. 4 provides the positive feature correspondence in our contrastive learning framework. Furthermore, we swap ^$\widehat{\mathcal{F}}^{t}$ when generating the target mean embedding features to form a symmetric representation. In this way, the correspondence is encouraged from both $t\to{t+1}$ and ${t+1}\to t$, which leads to the following optimization objective: ${\mathcal{L}}^{\text{tmp}}={\mathcal{L}}^{t\to{t+1}}+{\mathcal{L}}^{{t+1}\to t}$.
@@
Point to Segment Regularization. To pull close the LiDAR points belonging to the same instance at timestamp $t$, we minimize the distance between the point feature ^$\widehat{\mathcal{F}}^{t}$ and the corresponding mean cluster feature $\mathcal{C}^{t}$. To implement this, we leverage a max-pooling function to pool ^$\widehat{\mathcal{F}}^{t}$ according to the segments to obtain $\mathcal{C}^{t}=\{\mathcal{C}^{t}_{1},\mathcal{C}^{t}_{2},...,\mathcal{C}^{t}_{M_{k}}\}$, where $\mathcal{C}^{t}_{M_{k}}\in\mathbb{R}^{1\times D}$. The point-to-segment regularization is thus achieved via the ${\mathcal{L}}^{\text{p2s}}$ loss function as follows:
$-\frac{1}{M_{k}N_{k}}{\sum_{i=1}^{M_{k}}\sum_{a=1}^{N_{k}}\log\left[\frac{e^{\left(\left<{\mathbf{c}}^{t}_{i},{\widehat{\mathbf{f}}}^{t}_{i,a}\right>/\tau\right)}}{\sum_{j\neq i}e^{\left(\left<{\mathbf{c}}^{t}_{i},{\widehat{\mathbf{f}}}^{t}_{j,a}\right>/\tau\right)}+e^{\left(\left<{\mathbf{c}}^{t}_{i},{\widehat{\mathbf{f}}}^{t}_{i,a}\right>/\tau\right)}}\right]},$ (5)
@@
where $N_{k}$ represents the number of points within the corresponding segment. The final optimization objective is to minimize the aforementioned semantic spatial consistency loss $\mathcal{L}^{\text{vfm}}$, temporal consistency loss $\mathcal{L}^{\text{tmp}}$, and the point-to-segment regularization loss $\mathcal{L}^{\text{p2s}}$.

### v2

@@ Abstract
Recent advancements in vision foundation models (VFMs) have revolutionized visual perception in 2D, yet their potential for 3D scene understanding, particularly in autonomous driving applications, remains underexplored. In this paper, we introduce LargeAD, a versatile and scalable framework designed for large-scale 3D pretraining across diverse real-world driving datasets. Our framework leverages VFMs to extract semantically rich superpixels from 2D images, which are aligned with LiDAR point clouds to generate high-quality contrastive samples. This alignment facilitates cross-modal representation learning, enhancing the semantic consistency between 2D and 3D data. We introduce several key innovations: (i) VFM-driven superpixel generation for detailed semantic representation, (ii) a VFM-assisted contrastive learning strategy to align multimodal features, (iii) superpoint temporal consistency to maintain stable representations across time, and (iv) multi-source data pretraining to generalize across various LiDAR configurations. Our approach achieves substantial gains over state-of-the-art methods in linear probing and fine-tuning for LiDAR-based segmentation and object detection. Extensive experiments on 11 large-scale multi-sensor datasets highlight our superior performance, demonstrating adaptability, efficiency, and robustness in real-world autonomous driving scenarios.
@@ 1 Introduction
The emergence of large language models (LLMs) [1, 2, 3, 4] has transformed natural language processing, paving the way for similar breakthroughs in computer vision through vision foundation models (VFMs) such as SAM [5], X-Decoder [6], and SEEM [7]. These models have demonstrated remarkable capabilities in extracting rich pixel-level semantics from 2D images. However, extending these advancements to the 3D domain remains a largely unexplored frontier. As autonomous driving applications increasingly rely on 3D data from LiDAR sensors, there is a growing need to transfer the success of VFMs in 2D vision to 3D scene understanding.
Accurate perception of LiDAR point clouds is crucial for safe autonomous driving and advanced driver-assistance systems [8, 9, 10, 11, 12, 13]. Traditional LiDAR point cloud models often depend on large annotated datasets, which are costly and time-consuming to create [14, 15]. To alleviate this challenge, research has explored semi-supervised [16, 17] and weakly-supervised [15, 18] methods. However, these approaches suffer from limited generalizability, especially when faced with diverse sensor configurations, e.g., different LiDAR beam numbers, camera placements, sampling rates, and potential sensor corruptions [19, 20, 9, 21, 22, 23, 24, 25, 26, 27]. This limitation poses a significant challenge for real-world scalability.
@@
A core innovation in our framework is the use of VFMs to generate semantically enriched superpixels from camera images, which are then aligned with LiDAR data to construct high-quality contrastive samples, as shown in Fig. 1. These semantic superpixels provide enhanced 2D-3D correspondences that capture object-level coherence, reducing the errors commonly associated with over-segmentation and “self-conflict” in contrastive learning [29]. This alignment significantly improves performance in downstream tasks, including 3D object detection and segmentation.
@@
Figure 1: Comparisons between i) the conventional image-to-LiDAR data pretraining frameworks [30, 31, 29] and $i$) our proposed large-scale cross-sensor data pretraining (LargeAD). Our approach combines heterogeneous data sources for representation learning, which achieves superior robustness and scalability. Different from previous work, our framework encourages representation learning across different datasets, which largely enhances the generalizability.
@@
Figure 2: Illustration of image-to-LiDAR data pretraining using i) the heuristic SLIC algorithm [32] and $i$) different vision foundation models (VFMs). Images in the first row are the superpixels generated by different methods, where each color represents one distinct segment. The LiDAR point clouds from the second row are the superpoints grouped by projecting superpixels to 3D using camera-LiDAR correspondence. The third row shows the linear probing results after data pretraining.
@@ 2.1 LiDAR-Based Scene Understanding
For autonomous vehicles, accurate and dense 3D perception is crucial for safe navigation [10, 33, 34]. Researchers have developed various methods for point cloud segmentation, including those based on raw points [35, 36, 37, 38, 39], range views [40, 41, 42, 43, 44, 45, 46], bird’s eye views [47, 48], voxels [49, 50, 51, 52], and multi-view fusion [53, 54, 55, 56, 57, 58]. Despite achieving significant advancements, these models typically rely on extensively annotated datasets, posing scalability issues [19]. To alleviate the annotation burden, recent studies have explored semi-supervised [16, 17, 59], weakly-supervised [15, 60, 61, 62, 18], and active learning [63, 64, 65] methods, as well as domain adaptation techniques [66, 67, 68, 69, 70, 71]. This work utilizes a self-supervised learning strategy, distilling knowledge from VFMs through camera-to-LiDAR associations, thereby eliminating the need for manual annotations during pretraining.
@@ 2.3 Representation Learning in 3D
3D self-supervised learning has its roots in image-based techniques and typically focuses on object-centric point clouds [76, 77, 78, 79, 80] or indoor scenes [81, 82, 83, 84, 85] using pretext tasks [86, 87, 88, 89, 90], contrastive learning [91, 92, 93, 94, 95, 96, 97], or mask modeling [98, 99, 100, 101]. These methods often lack the necessary scale and diversity for outdoor driving scenes [102, 103, 104]. Efforts such as PointContrast [105], DepthContrast [106], and SegContrast [107] have pioneered contrastive objectives for small-scale point clouds. Recently, Sautier et al. [30] introduced SLidR, the first method for image-to-LiDAR representation distillation in cross-modal self-supervised learning on large-scale point clouds. Mahmoud et al. [31] further refined this approach with semantically tolerant contrastive constraints and a class-balancing loss. SuperFlow [108] introduced a spatiotemporal consistency framework to efficiently capture dynamic cues across multiple timestamps. Our framework builds upon SLidR [30] by leveraging VFMs [5, 6, 7] to create a more effective cross-modal contrastive objective. We also introduce a superpoint temporal consistency regularization to enhance feature learning and robustness in diverse and dynamic real-world driving scenarios.
@@ 2.4 Multi-Dataset Utilization
Leveraging multiple datasets has emerged as a promising approach for improving the generalization of LiDAR-based models in autonomous driving [109]. Recent works like MDT3D [110] and MS3D++ [111] explored multi-source training for 3D object detection, while addressing challenges such as label space conflicts. Similarly, methods like M3Net [112] utilized unified label spaces for semantic segmentation, demonstrating the advantages of multi-dataset learning. In this work, we extend these efforts by proposing a multi-source pretraining strategy that incorporates data from diverse LiDAR datasets, each with unique sensor setups and conditions [113, 114, 115, 12, 116]. By aligning 2D and 3D features across different sources, our approach improves cross-modal consistency and robustness. To the best of our knowledge, this is the first comprehensive study on multi-source pretraining for image-to-LiDAR representation learning, allowing our framework to generalize more effectively across varied driving environments and sensor configurations.
@@ 3 Image-to-LiDAR Data Pretraining
Figure 3: Overview of the VFM-driven image-to-LiDAR contrastive learning framework. Given a pair of LiDAR point cloud $\mathcal{P}^{t}$ and camera image $\mathcal{I}^{t}$ captured at timestamp $t$, along with another LiDAR point cloud $\mathcal{P}^{t+n}$ captured at timestamp $t+n$, we generate semantic superpixels using vision foundation models (VFMs). The corresponding superpoints are obtained by projecting image pixels onto the point cloud. Two key objectives are established: i) spatial contrastive learning between paired LiDAR and camera features, and $i$) temporal consistency regularization between point segments from $\mathcal{P}^{t}$ and $\mathcal{P}^{t+n}$.
@@ 3.1 Problem Formulation
Let us define a point cloud $\mathcal{P}=\{\mathbf{p}_{i},\mathbf{e}_{i}\mid i=1,\ldots,N\}$ consisting of $N$ points collected by a LiDAR sensor. Each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ denotes the 3D coordinates, while $\mathbf{e}_{i}\in\mathbb{R}^{L}$ represents its feature embedding, such as intensity, elongation, etc. This work aims to transfer knowledge from a set of surround-view images $\mathcal{I}=\{\mathbf{I}_{i}\mid i=1,\ldots,V\}$, captured by a total of $V$ synchronized RGB cameras, to the point cloud $\mathcal{P}$. Each image $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ has a spatial resolution defined by height $H$ and width $W$.
Given that the LiDAR and camera sensors are assumed to be well-calibrated [117, 118], each LiDAR point $\mathbf{p}_{i}=(x_{i},y_{i},z_{i})$ can be projected onto the image plane as a pixel $\hat{\mathbf{p}}_{i}=(u_{i},v_{i})$ using the following coordinate transformation:
@@ 3.3 Contrastive Learning Objective
Let $F_{\theta_{p}}:\mathbb{R}^{N\times(3+L)}\to\mathbb{R}^{N\times C}$ represent a LiDAR point cloud encoder with trainable parameters $\theta_{p}$, which processes a point cloud $\mathcal{P}$ and outputs a CC–dimensional feature for each point. Additionally, let $G_{\theta_{i}}:\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times E}$ be an image encoder with parameters $\theta_{i}$, initialized from 2D self-supervised pretrained models. To compute the superpixel-driven contrastive loss, we construct trainable projection heads $H_{\omega_{p}}$ and $H_{\omega_{i}}$ which map the 3D point features and 2D image features into the same $D$-dimensional embedding space. The point projection head $H_{\omega_{p}}:\mathbb{R}^{N\times C}\to\mathbb{R}^{N\times D}$ is a linear layer followed by $\ell_{2}$-normalization. The image projection head $H_{\omega_{i}}:\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times E}\to\mathbb{R}^{H\times W\times D}$ consists of a convolution layer with a $1\times 1$ kernel, followed by a fixed bi-linear interpolation layer in the spatial dimension, and outputs with $\ell_{2}$-normalization.
The goal is to distill the 2D network’s knowledge into the 3D network, ensuring that each semantic superpoint feature closely correlates with its corresponding semantic superpixel feature. Specifically, the superpixels $\Phi_{\mathcal{S}}$ and superpoints $\Phi_{\mathcal{O}}$ are used to group pixel and point embedding features, respectively. An average pooling operation is applied to the grouped pixel and point embeddings to obtain superpixel embedding features $\mathbf{Q}\in\mathbb{R}^{M\times D}$ and superpoint embedding features $\mathbf{K}\in\mathbb{R}^{M\times D}$. The contrastive loss $\mathcal{L}^{\mathrm{slic}}=\mathcal{L}\left(\mathbf{Q},\mathbf{K}\right)$ is then defined as follows:
$\mathcal{L}^{\mathrm{slic}}=-\frac{1}{M}{\sum_{i=1}^{M}\log\left[\frac{e^{\left(\left<\mathbf{q}_{i},\mathbf{k}_{i}\right>/\tau\right)}}{\sum_{j=1}^{M}e^{\left(\left<\mathbf{q}_{i},\mathbf{k}_{j}\right>/\tau\right)}}\right]},$ (2)
where $\left<\mathbf{q}_{i},\mathbf{k}_{j}\right$> represents the scalar product between superpoint and superpixel embedding features, measuring their similarity. $\tau$ is a temperature parameter used to scale the similarity scores.
@@ 4.2 Semantic Spatial Consistency Learning
To implement this framework, we use the same trainable LiDAR point cloud encoder $F_{\theta_{p}}$ and frozen image encoder $G_{\theta_{i}}$ as described earlier, extracting features from the LiDAR point cloud and 2D images, respectively. For the contrastive loss, we employ the projection heads $H_{\omega_{p}}$ and $H_{\omega_{i}}$ from Sec. 3.2, projecting both point and image features into a shared DD–dimensional embedding space. Unlike the low-level cues generated by SLIC, the superpixels produced by VFMs are enriched with semantic information, leading to more coherent and meaningful representations.
@@
$\mathcal{L}^{\mathrm{vfm}}=-\frac{1}{M_{v}}{\sum_{i=1}^{M_{v}}\log\left[\frac{e^{\left(\left<\hat{\mathbf{q}}_{i},\hat{\mathbf{k}}_{i}\right>/\tau\right)}}{\sum_{j=1}^{M_{v}}e^{\left(\left<\hat{\mathbf{q}}_{i},\hat{\mathbf{k}}_{j}\right>/\tau\right)}}\right]}.$ (3)
@@ 4.3 Instance Superpoint Temporal Consistency
Superpoint Temporal Consistency. We leverage the clustered segments to compute the temporal consistency loss among related semantic superpoints [^1] Here, we assume $n=1$, that is, the current and the next LiDAR frames, without loss of generalizability.. Specifically, given a sampled temporal pair $\widetilde{\mathcal{P}}^{t}$ and $\widetilde{\mathcal{P}}^{t+1}$ and their corresponding segments $\widetilde{\mathcal{K}}^{t}$ and $\widetilde{\mathcal{K}}^{t+1}$, we compute the point-wise features $\widehat{\mathcal{F}}^{t}\in\mathbb{R}^{N\times D}$ and $\widehat{\mathcal{F}}^{t+1}\in\mathbb{R}^{N\times D}$ from the point projection head $H_{\omega_{p}}$. As for the target embedding, we split the point features $\widehat{\mathcal{F}}^{t}$ and $\widehat{\mathcal{F}}^{t+1}$ into $M_{k}$ groups by segments $\widetilde{\mathcal{K}}^{t}$ and $\widetilde{\mathcal{K}}^{t+1}$. Then, we apply an average pooling operation on $\widehat{\mathcal{F}}^{t+1}$ to get $M_{k}$ target mean feature vectors $\widehat{\mathcal{F}}^{t+1}=\{\widehat{\mathcal{F}}^{t+1}_{1},\widehat{\mathcal{F}}^{t+1}_{2},...,\widehat{\mathcal{F}}^{t+1}_{M_{k}}\}$, where $\widehat{\mathcal{F}}^{t+1}_{M_{k}}\in\mathbb{R}^{1\times D}$. Let the split point feature $\widehat{\mathcal{F}}^{t}$ be $\widehat{\mathcal{F}}^{t}=\{\widehat{\mathcal{F}}^{t}_{1},\widehat{\mathcal{F}}^{t}_{2},...,\widehat{\mathcal{F}}^{t}_{M_{k}}\}$, where $\widehat{\mathcal{F}}^{t}_{M_{k}}\in\mathbb{R}^{k\times D}$ and $k$ is the number of points in the corresponding segment. We compute the temporal consistency loss ${\mathcal{L}}^{t\to{t+1}}$ to minimize the differences between the point features in the current frame (timestamp $t$) and the corresponding segment mean features from the next frame (timestamp $t+1$) as ${\mathcal{L}}^{t\to{t+1}}$ as follows:
$\mathcal{L}^{t\to{t+1}}=-\frac{1}{M_{k}}{\sum_{i=1}^{M_{k}}\log\left[\frac{e^{\left(\left<{\widehat{\mathbf{f}}}^{t}_{i},{\widehat{\mathbf{f}}}^{t+1}_{i}\right>/\tau\right)}}{\sum_{j=1}^{M_{k}}e^{\left(\left<{\widehat{\mathbf{f}}}^{t}_{i},{\widehat{\mathbf{f}}}^{t+1}_{j}\right>/\tau\right)}}\right]}.$ (4)
@@
Since the target embedding for all points within a segment in the current frame serves as the mean segment representation from the next frame, this loss will force points from a segment to converge to a mean representation while separating from other segments, implicitly clustering together points from the same instance. Fig. 4 provides the positive feature correspondence in our contrastive learning framework. Furthermore, we swap $\widehat{\mathcal{F}}^{t}$ when generating the target mean embedding features to form a symmetric representation. In this way, the correspondence is encouraged from both $t\to{t+1}$ and ${t+1}\to t$, which leads to the following optimization objective: ${\mathcal{L}}^{\mathrm{tmp}}={\mathcal{L}}^{t\to{t+1}}+{\mathcal{L}}^{{t+1}\to t}$.
@@
Point to Segment Regularization. To pull close the LiDAR points belonging to the same instance at timestamp $t$, we minimize the distance between the point feature $\widehat{\mathcal{F}}^{t}$ and the corresponding mean cluster feature $\mathcal{C}^{t}$. To implement this, we leverage a max-pooling function to pool $\widehat{\mathcal{F}}^{t}$ according to the segments to obtain $\mathcal{C}^{t}=\{\mathcal{C}^{t}_{1},\mathcal{C}^{t}_{2},...,\mathcal{C}^{t}_{M_{k}}\}$, where $\mathcal{C}^{t}_{M_{k}}\in\mathbb{R}^{1\times D}$. The point-to-segment regularization is thus achieved via the ${\mathcal{L}}^{\mathrm{p2s}}$ loss function as follows:
$\mathcal{L}^{\mathrm{p2s}}=-\frac{1}{M_{k}N_{k}}{\sum_{i=1}^{M_{k}}\sum_{a=1}^{N_{k}}\log\left[\frac{e^{\left(\left<{\mathbf{c}}^{t}_{i},{\widehat{\mathbf{f}}}^{t}_{i,a}\right>/\tau\right)}}{\sum_{j=1}^{M_{k}}e^{\left(\left<{\mathbf{c}}^{t}_{i},{\widehat{\mathbf{f}}}^{t}_{j,a}\right>/\tau\right)}}\right]},$ (5)
@@
where $N_{k}$ represents the number of points within the corresponding segment. The final optimization objective is to minimize the aforementioned semantic spatial consistency loss $\mathcal{L}^{\mathrm{vfm}}$, temporal consistency loss $\mathcal{L}^{\mathrm{tmp}}$, and the point-to-segment regularization loss $\mathcal{L}^{\mathrm{p2s}}$.